The table below shows the number of male and female students enrolled in nursing at a particular university for a recent semester 

a) Find the probability that a randomly selected student is male, given that the student is a nursing major 

(b) Find the probability that a randomly selected student is a nursing major, given that the student is male. 

$$ \begin{array}{cccc}  & \text { Nursing Majors } & \text { Non-nursing majors } & \text { Total } \\ \text { Males } & 95 & 1086 & 1181 \\ \text { Females } & 724 & 1691 & 2415 \\ \text { Total } & 819 & 2777 & 3596 \end{array} $$

(a) Find the probability that a randomly selected student is male, given that the student is a nursing major The probability is _______ (Round to three decimal places as needed.) 

(b) Find the probability that a randomly selected student is a nursing major, given that the student is male. The probability is _______ (Round to three decimal places as needed.)
Answer #1
Answer #2

SOLUTION :


a.


P(Male | Nursing majors) = 95/819 = 0.116 (ANSWER)

(We select a male from the Nursing majors space)


b.


P(Nursing majors | Male) = 95/1181 = 0.080 (ANSWER)

(We select a Nursing majors in Male space)




answered by: Tulsiram Garg